# Chain-of-Translation Prompting (CoTR): A Novel Prompting Technique for Low Resource Languages

### v1

@@ 3.1 Chain of Translation Prompting
    Comparison with Direct Prompting: To evaluate the effectiveness of the Chain of Translation Prompting, we compare its results against the traditional method of directly prompting the models to process the Marathi text without performing translation.
@@
Figure 4: Classification Task using Chain of Translation Prompting
@@ 3.3 Evaluation Methodology
Table 3: Error percentage in the classification task across 5 models (these are the weighted averages and the numbers are percentages)
Model Dataset Standard Prompt CoTR Prompt Average Standard Prompt Average CoTR Prompt
gpt-4o MahaSent-GT 20.38 18.44 13.57 11.25
MahaNews-SHC 3.06 2.04
MahaHate 16.83 12.87
gpt-4o mini MahaSent-GT 20.38 19.41 20.19 15.23
MahaNews-SHC 6.12 4.08
MahaHate 33.66 21.78
llama3-405b MahaSent-GT 31.06 27.18 19.86 16.22
MahaNews-SHC 7.14 6.12
MahaHate 20.89 14.85
llama3-8b MahaSent-GT 35.92 27.18 29.13 23.84
MahaNews-SHC 10.20 7.14
MahaHate 40.59 36.63
gemma9b MahaSent-GT 33.98 27.18 22.18 22.51
MahaNews-SHC 10.20 10.20
MahaHate 21.78 29.70
@@ 4.2 Generation Task
Table 4: Rouge-L score for 3 approaches on the headline generation task

### v2

@@ 3.1 Chain of Translation Prompting
Figure 4: Classification Task using Chain of Translation Prompting
@@
    Direct Prompting: To evaluate the effectiveness of the Chain of Translation Prompting, we compare its results against the traditional method of directly prompting the models to process the Marathi text without performing translation.
@@
    Google Translate + Prompting: In this approach, Marathi sentences were translated into English using Google Translate. The translated English sentences, along with English prompts, were then used by LLMs to perform the desired classification tasks. This method represents a straightforward "translate-and-test" approach, serving as a baseline for comparison.
  6. 6.
@@ 3.3 Evaluation Methodology
Table 3: Error percentage in the classification task across 5 models (these are the weighted averages and the numbers are percentages). Standard Prompt - Prompt the LLM to perform the task using the given Marathi context. CoTR Prompt - Prompt the LLM to translate the Marathi context to English and then perform the task. Google Translate - Translate the Marathi context to English using Google Translate and then prompt the LLM to perform the task in English.
Model Dataset Standard Prompt CoTR Prompt Google Translate Average Standard Prompt Average CoTR Prompt Avg Google Translate Prompt
gpt-4o MahaSent 20.38 18.44 25.00 13.57 11.25 18.70
MahaNews 3.06 2.04 6.12
MahaHate 16.83 12.87 26.70
gpt-4o mini MahaSent 20.38 19.41 33.00 20.19 15.23 24.40
MahaNews 6.12 4.08 9.18
MahaHate 33.66 21.78 30.70
llama3-405b MahaSent 31.06 27.18 22.00 19.86 16.22 18.70
MahaNews 7.14 6.12 6.12
MahaHate 20.89 14.85 27.72
llama3-8b MahaSent 35.92 27.18 30.00 29.13 23.84 24.00
MahaNews 10.20 7.14 9.18
MahaHate 40.59 36.63 32.60
gemma9b MahaSent 33.98 27.18 29.00 22.18 22.51 22.40
MahaNews 10.20 10.20 11.20
MahaHate 21.78 29.70 26.70
@@ 4.1 Classification Task
The results from the CoTR approach show significant improvement over the standard Google Translate method as well. We manually reviewed the translated sentences and found out that translations by large language models (LLMs), such as GPT-4 and GPT-4 Mini, captured meanings and nuances more effectively than Google Translate. While LLMs conveyed the intended meaning with subtlety, Google Translate produced literal translations, which sometimes failed to capture the full sense of the sentences.
For GPT-4 and GPT-4 Mini, the direct translation approach surpassed Google Translate’s performance, as the nuances of some of the sentences did not get extracted completely by the google translator. As GPT-4 and GPT-4 Mini are stronger models the direct prompting is working better than Google translator approach.
@@
Table 4: Sample examples for the headline generation task using CoTR